OpenStudy (anonymous):

The mean, mode, median, range of 180 170 190 200 130 30 180 160

OpenStudy (anonymous):

mean = 155 mode = 130 median = 180 + 170 /. 2 = 175 range = 200 - 30 = 170

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

plz click good answer

OpenStudy (anonymous):

therrr u go